OpenStudy (anonymous):

-3(x+7)^2=-42. can anyone show me how to solve this problem and if so can you please show work or give me a brief description?

OpenStudy (anonymous):

divide both sides by \(-3\) first to get rid of the \(-3\) out front what do you get?

OpenStudy (anonymous):

(x+7)^2=14?

OpenStudy (anonymous):

looks good to me then take the square root of both sides, don't forget the \(\pm\)

OpenStudy (anonymous):

in other words, just write \[x+7=\pm\sqrt{14}\]

OpenStudy (anonymous):

then solve by subtracting \(7\)

OpenStudy (anonymous):

subtract 7 from radical 14? like this 7-\[\sqrt{14}\]

OpenStudy (anonymous):

no, subtract 7 \[-7\pm\sqrt{14}\]

OpenStudy (anonymous):

ok, thank you(: